1. Find the weak form for the following boundary value problem: (Uxx+u = 0 u(0) = 0 (uz (1) = -1
Answer #1

given that Bernary conda Yoxy tuzo 460)=0 2 41x (1) = -1 dy +420 62+1) 4 = 0 (m2+1) 4 = 0 2 t =0 m2 -1, ima I i 4 * A cos 1aFirst convert eqaution into second order differential equation,

then find value of m

after that find CF of equation in which have two constant A and B

The constant are solved by applying boundry conditions,

so finally putting the value of constants we get final solution